Verify rolles theorem for f(x)=2x²+x²-4x-2 in (-√2,√2)

Grade:12th pass

1 Answers

Subham Patel
4 years ago
This polynomial equation  
f(X) = 2x^2 +x^2 -4x -2 does not verify rolles theorem bcz its third conditions f(a)=f(b) is not satisfied 
On solving we get f(a )= 4+4(2)^1/2 
Whereas f(b) = 4-4(2)^1/2 , hence f(a)  is unequal to f(b) . Thus roll es theorem is not satisfied …
                                                                Thank u .....